OpenStudy (anonymous):

What is 5/6 divided by 1/3?

OpenStudy (anonymous):

\[\frac{ 5 }{ 6 } \div \frac{ 1 }{ 3 }\] is the same as \[\frac{ 5 }{ 6 } * \frac{ 3 }{ 1 }\] \[\frac{ 5 }{ 2 } * \frac{ 1 }{ 1 }\] \[\frac{ 5 }{ 2 }\] 2.5

OpenStudy (anonymous):

Ok Thank you